OpenStudy (anonymous):

C(x) = 4x2; R(x) = x3 + 5x + 10 Find all values of x for which the marginal profit is zero. Interpret your answer.

OpenStudy (anonymous):

Assuming C(x) and R(x) are total cost/total revenue (respectively).... Take the derivative of each to find it's slope (the change per each unit). C'(x) = 8x R'(x) = 3x^2 + 5, or MC = 8x and MR = 3x^2+5 MP (marginal profit) = MR - MC. MP = 3x^2 - 8x + 5. Factor that to find the roots, and set each to zero. 3x^2-5x-3x+5 = 0 x(3x-5) - 1(3x-5) = 0 (x-1) (3x-5) = 0 MP = 0 @ x = 1 and 5/3